Required information Camp Rainbow offers overnight summer camp programs for children ages 10-14 every summer during June andRequired: 1. Perform a least-squares regression analysis on Camp Rainbows data. 2. Using the regression output, create a cosRequired 1 Required 2 Required 3 Using the regression output, create a cost equation (y = a + bx) for estimating Camp RainbowRequired 1 Required 2 Required 3 Using the least-squares regression results, calculate the camps expected operating cost if
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Answer 1
Least Squares Regression
The equation of a line is y=mx+b
X= number of copies
y=total cost
m= (N Σ(xy) − Σx Σy)/ (N Σ(x2) − (Σx)^2)
Week Campers (x) Costs (y) x2 xy
1              188.00             13,340.00           35,344.00       2,507,920.00
2              140.00               9,940.00           19,600.00       1,391,600.00
3              174.00             13,780.00           30,276.00       2,397,720.00
4              152.00             10,130.00           23,104.00       1,539,760.00
5              180.00             11,210.00           32,400.00       2,017,800.00
6              192.00             11,210.00           36,864.00       2,152,320.00
7              102.00               9,590.00           10,404.00           978,180.00
8              114.00               8,000.00           12,996.00           912,000.00
Total          1,242.00             87,200.00        200,988.00     13,897,300.00
N= number of weeks= 8.
m= (N Σ(xy) − Σx Σy)
(N Σ(x2) − (Σx)2)
m= 8*13897300- 1242*87200
8*200988- 1242^2
m= 111178400- 108302400
1607904-1542564
m= 2,876,000.00
       65,340.00
m=                44.02
Variable 1                44.02
b= (Σy − m Σx)/N
b= 87200- 44.02*1242
                 8.00
b= 87200- 54667.7685950413
                 8.00
b=        32,532.23
                 8.00
b=          4,066.53
Intercept          4,066.53
Answer 2
The equation of a line is y=mx+b
y= 44.02x+4066.53
Total cost= $ 4,066.53+ $ 44.02* Number of campers
Answer 3
Cost of 102 campers copies 44.02*102+4066.53
Cost of 102 campers copies $      8,556.15